OpenStudy (anonymous):

HELP ME PLEASE the function f is defined as follows f(x)=4+x if x < 0 x^2 if x >or equal to 0 A) find the domain of the function B) Locate any intercepts C) Graph the function D) based on the graph, find the range E) is f continuous on its domain?

OpenStudy (blurbendy):

f(x) = x^2 if x >= to 0?

OpenStudy (anonymous):

yes

OpenStudy (blurbendy):

The domain of the first one would be all negative numbers to infinity, because x has to be < 0. Can you figure out the domain of f(x) = x^2 if x >= 0 ?

OpenStudy (anonymous):

wouldn't that be (0, infinity) so would the domain be (-infinity, infinity)

OpenStudy (blurbendy):

it would need to be [0, infinity), you need the bracket because x is greater than or EQUAL to 0. if it can equal 0, then you have to include it in the domain, and you do that by using [ sidenote: infinity never has has a [

OpenStudy (anonymous):

so the domain would be [0, infinity) what about the other equation?

OpenStudy (blurbendy):

the domain of f(x) = x^2 is [0, infinity)

OpenStudy (anonymous):

yes but the overall domain would be (-infinity, infinity) right?

OpenStudy (blurbendy):

the domain of the first one is (0, -infinity)

OpenStudy (blurbendy):

you dont include 0 this time

OpenStudy (anonymous):

so the overall equation for both domain would be (-infinity, 0) U [0, infinity) ??

OpenStudy (blurbendy):

yes!!

OpenStudy (anonymous):

then what would be the range???

OpenStudy (blurbendy):

so what would happen if you started plugging in negative numbers for x in the first equation? what's the first answer you would get

OpenStudy (anonymous):

the out come would be neg.

OpenStudy (blurbendy):

4 + (-1) = 3 4 + (-2) = 2 4 + (-infinity) = -infinity so what's the range?

OpenStudy (anonymous):

range would be (-infinity, infinity)...??

OpenStudy (blurbendy):

(-infinity, 3] for the first one

OpenStudy (blurbendy):

what about the second?

OpenStudy (anonymous):

wouldn't it be (-infinity , 4) second would be [o, infinity)

OpenStudy (blurbendy):

you can write it either way

OpenStudy (blurbendy):

and yes, that is the second one

OpenStudy (anonymous):

then the interception would be (0,4) and [0,0]

OpenStudy (blurbendy):

have you graphed the function?

OpenStudy (anonymous):

yes the first equation y intercept is 4 but its an open circle, also have a x intercept of [-4,0] and the second equation stats at (0,0) and continues to infinity for got about the x intercept on the first intercept

OpenStudy (blurbendy):

(0,0) and (1,1) f(-4) = 4 + (-4) = 0 f(0) = (0)^2 = 0 f(3) = 4 + (-3) = 1 f(1) = x^2 = 1

OpenStudy (anonymous):

so would it be [0,0] and [-4,0] then why (1,1)

OpenStudy (blurbendy):

intercepts are ordered pairs, you write them like this ( x , y ) the only places they intersect are (0,0) and (1,1)

OpenStudy (blurbendy):

look at why that is two posts up

OpenStudy (anonymous):

but they don't intercept at (0,0) or (1,1) on the grpah

OpenStudy (blurbendy):

i just graphed it in my calculator. they dont intersect at whole numbers actually...

OpenStudy (blurbendy):

wait, there are constraints

OpenStudy (blurbendy):

there arent any intercepts.

OpenStudy (blurbendy):

because the functions are constrained

OpenStudy (anonymous):

i think the intercept would be any intercept on the y axis and the x axis and yes there are constrains next to the equation would the constrained

OpenStudy (blurbendy):

intercepts are where the two equations equal the same thing given the SAME input. This can never occur because the functions are constrained

OpenStudy (anonymous):

 is somewhat how the graph looks like